OpenStudy (mandineric):

The functions f(x) and g(x) are described below: f(x) = 7x + 9 g(x) = 7x − 4 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (amorfide):

let's say you have f(x)=ax+b b is a translation in the positive y direction, of b units a is a stretch factor of a, parallel to the y axis f(x)=ax-b -b is a translation in the negative y direction of b units a is a stretch factor of a, parallel to the y axis so f(x)=3x+9 would be translated 9 units in the positive y direction and stretched by a factor of 3 in the y axis so if I had 2 functions f(x)=3x+9 g(x)=3x+6 I have the same stretch factor, but different translations so the translation is the difference of 9 and 6 9-6=3 hence it must be translated 3 units since im going from 9 to 6, im going down, so i will translate it in the negative y directions hence i must translate my graph 3 units in the negative y direction sooooo go back to your question if you have f(x)=7x+9 g(x)=7x-4 what did you translate by to get from 9 to -4

OpenStudy (mandineric):

can you show me how to do it

OpenStudy (mandineric):

hi every one on the internet